A production process is sampled 24 times with a sample size of 6 yields, an overall mean of 25.2, and an average range of 0.8. Determine the control limits for the x-bar and R-charts.
Answer #1

For the x-chart, A2 = 0.483; thus the control limits are

UCL = 25.2 + 0.483(0.8) = 25.5864

LCL = 25.2 - 0.483(0.8) = 24.8136

For the R-chart,

D4= 2.004; D3= 0;

the control limits are

UCL = D4*R-bar = 2.004(0.8) = 1.6032

LCL = D3*R-bar = 0